# Taxonomy and Estimation of Multiple Breakpoints in High-Dimensional Factor Models

### v2

@@ 2.2 Estimator
Let $\hat{G}=(\hat{g}_{1},...,\hat{g}_{T})^{\prime}$ denote the full-sample PCA estimator for $G$ subject to the normalization condition that $G^{{}^{\prime}}G/T=I_{r}$ and $\Lambda^{{}^{\prime}}\Lambda$ being diagonal. The number of pseudo-factors can be estimated using Bai and Ng (2002).  [^1] Other methods for determining the number of factors include the Eigenvalue Ratio (ER) and Growth Ratio (GR) (Ahn and Horenstein, 2013), the empirical distribution estimator (Onatski, 2010), and the bridge estimator (Caner and Han, 2014). For a given set of splitting points $(k_{1},\cdots,k_{m_{0}})$, the QML objective function is expressed as:
@@
where $K_{\eta}=\{(k_{1},\cdots,k_{m_{0}}):k_{j}-k_{j-1}\geq T\eta,1<k_{j}<T,j=1,\cdots,m_{0}+1\}$ with $\eta\in(0,1)$.
Remark 1
Bai and Perron (2003) introduced a dynamic programming algorithm that efficiently minimizes the sum of squared residuals by comparing all possible segments, requiring at most $O(T^{2})$ least-squares operations regardless of the number of breaks. In this paper, we adapt this approach to efficiently optimize the QML objective function over all possible $(k_{1},\cdots,k_{m_{0}})$.
@@
3 Type of breakpoints
The representation in (17) is convenient for theoretical analysis because one can control break types by setting $B_{j}$ for $j=1,\cdots,m_{0}+1$. To further illustrate the generality of model (17), we present two types of breakpoint structures within the combined regime $k_{j-1}^{0}+1,k_{j+1}^{0}]$ by varying the ranks of $B_{j}$ and $B_{j+1}$. Define the number of effective factors in the combined time interval $k_{j-1}^{0}+1,k_{j+1}^{0}]$ (combined regime) as  [^2] This is because $r_{j,j+1}$ $=\text{rank}([B_{j}F_{j}^{\prime},B_{j+1}F_{j+1}^{\prime}]^{\prime})=\text{rank}([B_{j}F_{j}^{\prime},B_{j+1}F_{j+1}^{\prime}][B_{j}F_{j}^{\prime},B_{j+1}F_{j+1}^{\prime}]^{\prime})$ $=\text{rank}([B_{j},B_{j+1}]diag(F_{j}^{\prime}F_{j},\ F_{j+1}^{\prime}F_{j+1})[B_{j},B_{j+1}]^{\prime})=\text{rank}([B_{j},B_{j+1}]).$
@@ 3 Type of breakpoints
Next, we will show there are two types of breakpoints.
@@
If $r_{j,j+1}>\max\{r_{j},r_{j+1}\}$, the dimension of the factor space over the combined regime $k_{j-1}^{0}+1,k_{j+1}^{0}]$ increases due to the structural break. This occurs when the factor loading matrices from regime $k_{j-1}^{0}+1,k_{j}^{0}]$ and regime $k_{j}^{0}+1,k_{j+1}^{0}]$ jointly span a larger space than either individual matrix. This space augmentation is analogous to a regression where a break in coefficients increases the number of effective regressors; specifically, the interaction of a break dummy with original regressors leads to an expanded regressor set. To illustrate how our framework accommodates this scenario, consider Data Generating Process (DGP) 1.A in Section 5. In DGP 1.A, the setting involves three regimes, each with $r_{j}=3$ effective factors. Since the loading coefficients for these three regimes are independently generated, the total number of pseudo-factors in the full sample is nine ($r=9$), resulting in $\Lambda$ being an $N\times 9$ matrix. Here, $B_{1}$, $B_{2}$, and $B_{3}$ are diagonal matrices of rank 3 such that $\Lambda B_{1}$, $\Lambda B_{2}$, and $\Lambda B_{3}$ collect the first three, middle three, and last three columns of $\Lambda$, respectively, representing the loadings on the effective factors for each regime.
If $r_{j,j+1}=\max\{r_{j},r_{j+1}\}$, we observe emerging or disappearing factors. For instance, when $r_{j,j+1}=r_{j}$ and $r_{j+1}<r_{j,j+1},$ some factors disappear after the breakpoint $k_{j}^{0}$, as illustrated by DGP 1.C.
Type 2. $r_{j,j+1}=\min\{r_{j},r_{j+1}\}$. In this case, the number of pseudo-factors in the combined regime $k_{j-1}^{0}+1,k_{j+1}^{0}]$ is the same as the number of factors in each regime, and we refer to this as a rotational change. Specifically, there exists a nonsingular $r\times r$ matrix $\bf R$ such that
@@
which satisfies $r_{j,j+1}=r_{j}=r_{j+1}=rank(B_{j})$. When there is one break ($m_{0}=1$) and $B_{1}$ is nonsingular, it is analogous to the rotational change for single break setup. For example, in DGP 1.D, setting $B_{2}=2B_{1}$ means that the factors after the breakpoint have twice standard deviation as those before the breakpoint, and hence $r_{1,2}=r_{1}=r_{2}$. Such a breakpoint suggests that a model exhibiting a rotational change in the factor covariance matrix, with unchanged loadings, can be equivalently regarded as a Type 2 change.
@@
Notably, the analysis of multiple breakpoints is considerably more complicated than for a single breakpoint. In a multiple breakpoint setting, $r_{j,j+1}$ can be less than the total number of pseudo-factors $r$. Specifically, it is possible that both $B_{j}$ and $B_{j+1}$ are singular, yet the $j$-th breakpoint is rotational. Consequently, relying solely on the singularity of $B_{j}$ (like the DBH rule) is insufficient for determining the breakpoint type. To understand this phenomenon, consider Equations (19) and (26): when $B_{j}$ is singular (and consequently $B_{j+1}$ is singular), the pseudo-factors in regimes $j$ and $j+1$ (i.e., $F_{j}B_{j}^{\prime}$ and $F_{j+1}B_{j+1}^{\prime}$) are not of full column rank. However, if we focus solely on the data in the combined interval $k_{j-1}^{0}+1,k_{j+1}^{0}]$ and fit a factor model for observations within this interval, the pseudo-factors in this model are of full column rank $r_{j,j+1}$. In contrast, when considering the full sample for t=1,,Tt=1,...,T, the pseudo-factors in the interval $k_{j-1}^{0}+1,k_{j+1}^{0}]$ become rank-deficient. Therefore, this singularity is not directly attributable to the $j$-th breakpoint itself but rather results from the expansion of the factor loadings space caused by other breakpoints, which renders the pseudo-factors in $k_{j-1}^{0}+1,k_{j+1}^{0}]$ rank-deficient. This highlights that for rotational changes, the singularity of $B_{j}$ and $B_{j+1}$ is not introduced by the $j$-th breakpoint but rather by the cumulative effect of structural changes at multiple locations. Therefore, relying solely on the singularity of $B_{j}$ and $B_{j+1}$ is insufficient to distinguish between type 1 and type 2 changes at the $j$-th breakpoint, as both types may be associated with singular $B_{j}$ and $B_{j+1}$. Situations involving rank deficiency for regimes before and after a rotational change occur only when multiple breakpoints are present; such scenarios do not arise in single-breakpoint cases, highlighting the complexity introduced by the multi-breakpoint setting.
The preceding discussion naturally raises the question of how to distinguish between the two types of changes in multi-breakpoint settings, particularly when the singularity of $B_{j}$ and $B_{j+1}$ (the DBH rule in a single-break setting) is no longer sufficient for distinction. The answer is to use the definition $r_{j}=r_{j+1}=r_{j,j+1}$ to determine if the $j$-th breakpoint is rotational. Evidently, if Equation (19) holds, then $r_{j}=r_{j+1}=r_{j,j+1}$ must also hold, implying that the $j$-th breakpoint is rotational. Conversely, another natural question arises: does $r_{j}=r_{j+1}=r_{j,j+1}$ always imply the existence of a nonsingular $R$ matrix satisfying Equation (19)? This is affirmed by Proposition 1, which demonstrates that Equation (19) is a necessary and sufficient condition for defining a Type 2 break. This proposition also explains why Type 2 is named a “rotational change”.
@@
Proposition 1 can equivalently be written as: There does not exist a nonsingular $\bf R$} such that $B_{j}{\bf R}=B_{j+1}$ $\Leftrightarrow$ $r_{j,j+1}>\min(r_{j},r_{j+1})$, which is corresponding to a type 1 change. To further illustrate this, consider two distinct cases for a setting involving two breakpoints. Case 1 involves three regimes characterized by $B_{1}$, $B_{2}$ and $B_{3}$; Case 2 involves $B_{1}$, $\tilde{B}_{2}$ and $B_{3}$. These matrices are defined as:
@@
But such nonsingular $\bf R$} does not exist for $\tilde{B}_{2}$ because the second row of the product $B_{1}{\bf R}$ is always zeros, which is not equal to that of $\tilde{B}_{2}$, i.e., for any $\bf R$},
@@
For Type 1 breakpoints, we will show that the QML estimators converge to the true breakpoints with probability approaching 1. Conversely, for Type 2 breakpoints, the difference between QML estimators and the true change points is stochastically bounded. The specific asymptotic theory is detailed in Section 4.
@@ 4.1 Assumptions
(i) $E\left\|f_{t}\right\|^{4}<M<\infty$, $E(f_{t}f_{t}^{{}^{\prime}})=\Sigma_{F}$, where $M$ is a positive constant, $\Sigma_{F}$ is positive definite, and $\frac{1}{k_{j}^{0}-k_{j-1}^{0}}\sum_{t=k_{j-1}^{0}+1}^{k_{j}^{0}}f_{t}f_{t}^{{}^{\prime}}\xrightarrow{p}\Sigma_{F}$ for $j=1,\cdots,m_{0}+1$;
@@
(iii) $B_{1},\cdots,B_{m_{0}}]$ is of full row rank.
@@
$\left\|\lambda_{i,j}\right\|\leq\bar{\lambda}<\infty$ for $j=1,\cdots,m_{0}+1$, $i=1,\cdots,N$, and $\left\|\frac{1}{N}\Lambda^{{}^{\prime}}\Lambda-\Sigma_{\Lambda}\right\|\rightarrow 0$ for some $r\times r$ positive definite matrix $\Sigma_{\Lambda}$.
@@
    For every $(i,j)$ and $\ell=1,\cdots,m_{0}+1$, $E\left|(k_{\ell}^{0}-k_{\ell-1}^{0})^{-1/2}\sum_{t=k_{\ell-1}^{0}+1}^{\ell_{j}^{0}}(e_{it}e_{jt}-E[e_{it}e_{jt}])\right|^{4}\leq M$.
@@
Define $\epsilon_{t}=f_{t}f_{t}^{{}^{\prime}}-\Sigma_{F}$. The Hájek-Rényi inequality applies to the processes $\{\epsilon_{t},t=k_{j-1}^{0}+1,\cdots,k_{j}^{0}\}$ and $\{\epsilon_{t},t=k_{j}^{0},\cdots,k_{j-1}^{0}+1\}$ for $j=1,\cdots,m_{0}+1$.
@@
There exists an $M<\infty$ such that
(i) for each $h=1,\cdots,T$:
$E\left(\max_{\begin{subarray}{c}k-s>(\log T)^{2}\end{subarray}}\frac{1}{k-s}\sum_{t=s+1}^{k}\left|\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\left[e_{ih}e_{it}-E(e_{ih}e_{it})\right]\right|^{2}\right)$ $\leq$ $M;$
(ii) $E\left(\max_{\begin{subarray}{c}k-s>(\log T)^{2}\end{subarray}}\frac{1}{k-s}\sum_{t=s+1}^{k}\left\|\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\lambda_{i}e_{it}\right\|^{2}\right)\leq M.$

### v3

@@ 2.2 Estimator
Let $\hat{G}=(\hat{g}_{1},...,\hat{g}_{T})^{\prime}$ denote the full-sample PCA estimator for $G$ subject to the normalization condition that $G^{{}^{\prime}}G/T=I_{r}$ and $\Lambda^{{}^{\prime}}\Lambda$ being diagonal. The number of pseudo-factors can be estimated following Bai and Ng (2002).  [^1] Other methods for determining the number of factors include the eigenvalue ratio and growth ratio (Ahn and Horenstein, 2013), a similar method using adjacent eigenvalue ratios proposed by Lam and Yao (2012), the empirical distribution estimator (Onatski, 2010), and the bridge estimator (Caner and Han, 2014). For a given set of splitting points $(k_{1},\cdots,k_{m_{0}})$, the QML objective function is expressed as follows:
@@
where $K_{\eta}=\{(k_{1},\cdots,k_{m_{0}}):0=k_{0}<k_{1}<\cdots<k_{m_{0}}<k_{m_{0}+1}=T,k_{j}-k_{j-1}\geq T\eta,j=1,\cdots,m_{0}+1\}$ and $\eta\in(0,1)$ is a small positive constant.
Remark 1 (Dynamic programming implementation for the QML estimator)
We explain how to compute the QML estimator in (4) using a dynamic programming (DP) algorithm in the spirit of Bai and Perron (1998, 2003). For any integers $0\leq s<e\leq T$, define
$\hat{\Sigma}(s,e)=\frac{1}{e-s}\sum_{t=s+1}^{e}\hat{g}_{t}\hat{g}_{t}^{\prime},\qquad\mathcal{Q}(s,e)=(e-s)\log\big|\hat{\Sigma}(s,e)\big|.$
Then for any partition $0=k_{0}<k_{1}<\cdots<k_{m}<k_{m+1}=T$, the QML objective satisfies the additivity property
$U_{NT}(k_{1},\ldots,k_{m})=\sum_{\ell=1}^{m+1}\mathcal{Q}(k_{\ell-1},k_{\ell}).$
Let $h=\lceil T\eta\rceil$ be the minimum segment length implied by the constraint set $K_{\eta}$. Define $D(m,e)$ as the minimum value of $\sum\mathcal{Q}(\cdot,\cdot)$ over all partitions of the first e observations $\{1,\ldots,e\}$ with exactly $m$ breaks such that each segment has length at least $h$. The DP recursion is
$D(0,e)=\mathcal{Q}(0,e),\qquad e\geq h,$
and for $m\geq 1$ and $e\geq(m+1)h$,
$D(m,e)=\min_{j\in\{mh,\ldots,e-h\}}\Big\{D(m-1,j)+\mathcal{Q}(j,e)\Big\}.$
For a given $m$, evaluating $D(m,e)$ for all $e\geq(m+1)h$ requires $O(T^{2})$ operations by varying the endpoint $e$ and the candidate breakpoint $j\leq e-h$. The final breakpoint sequence is recovered via backtracking using the stored argmin indices: let $\hat{k}_{m_{0}+1}=T$, and for $\ell=m_{0},\ldots,1$ set $\hat{k}_{\ell}$ equal to the minimizer of $D(\ell-1,j)+\mathcal{Q}(j,\hat{k}_{\ell+1})$ over $j\in\{\ell h,\ldots,\hat{k}_{\ell+1}-h\}$.
Algorithm 1 DP computation of $(\hat{k}_{1},\ldots,\hat{k}_{m_{0}})$ for a fixed $m_{0}$
1:Input: $\{\hat{g}_{t}\}_{t=1}^{T}$, $\eta\in(0,1)$, $h=\lceil T\eta\rceil$, a fixed number of breaks $m_{0}$.
2:Precompute prefix sums: $S_{0}=0_{r\times r}$ and $S_{t}=\sum_{u=1}^{t}\hat{g}_{u}\hat{g}_{u}^{\prime}$ for $t=1,\ldots,T$.
3:Segment costs: For all $0\leq s<e\leq T$ with $e-s\geq h$, set $\hat{\Sigma}(s,e)=(S_{e}-S_{s})/(e-s)$ and $\mathcal{Q}(s,e)=(e-s)\log|\hat{\Sigma}(s,e)|$.
4:DP initialization: For $e=h,\ldots,T$, set $D(0,e)=\mathcal{Q}(0,e)$.
5:for $m=1,\ldots,m_{0}$ do
6:  for $e=(m+1)h,\ldots,T$ do
7:   $D(m,e)=\min\limits_{j=mh,\ldots,e-h}\{D(m-1,j)+\mathcal{Q}(j,e)\}$.
8:   Store an argmin value: $A(m,e)=\arg\min\limits_{j=mh,\ldots,e-h}\{D(m-1,j)+\mathcal{Q}(j,e)\}$.
9:  end for
10:end for
11:Backtracking: Set $\hat{k}_{m_{0}+1}=T$ and $\hat{k}_{\ell}=A(\ell,\hat{k}_{\ell+1})$ for $\ell=m_{0},\ldots,1$.
12:Output: $(\hat{k}_{1},\ldots,\hat{k}_{m_{0}})$.
The estimator can be viewed as a two-step procedure: first estimate $\hat{G}$ by full-sample PCA, and then locate the breakpoints by minimizing the QML objective function. Unlike classical methods, our log-determinant objective exploits the near-singularity of $\hat{\Sigma}(k_{\ell-1},k_{\ell})$ when the segment boundaries approach true Type 1 breakpoints. This sensitivity to the rank-deficient structure enables the faster convergence rates established in Theorem 2, outperforming existing methods such as Barigozzi et al. (2018, 2025), Baltagi et al. (2017, 2021).
Once the breakpoints are estimated, one can apply split-sample PCA to consistently estimate the loadings and factors for each regime. Using data within the interval $\hat{k}_{j-1}+1,\hat{k}_{j}]$, one can recover the regime-specific $r_{j}$-dimensional factors $\hat{f}_{t}$ and loadings $\hat{\Lambda}_{j}$, subject to a local $r_{j}\times r_{j}$ rotation $\hat{H}_{j}$.
@@
3 Types of breakpoints
The representation in (17) is convenient for theoretical analysis because one can control break types by setting $B_{j}$ for $j=1,\cdots,m_{0}+1$. To further illustrate the generality of model (17), we present two types of breakpoint structures within the combined time interval $k_{j-1}^{0}+1,k_{j+1}^{0}]$ (combined regime) by varying the ranks of $B_{j}$ and $B_{j+1}$. We define the number of pseudo-factors in the combined regime $k_{j-1}^{0}+1,k_{j+1}^{0}]$ as  [^2] This is because $r_{j,j+1}$ $=\text{rank}([B_{j}F_{j}^{\prime},B_{j+1}F_{j+1}^{\prime}]^{\prime})=\text{rank}([B_{j}F_{j}^{\prime},B_{j+1}F_{j+1}^{\prime}][B_{j}F_{j}^{\prime},B_{j+1}F_{j+1}^{\prime}]^{\prime})$ $=\text{rank}([B_{j},B_{j+1}]diag(F_{j}^{\prime}F_{j},\ F_{j+1}^{\prime}F_{j+1})[B_{j},B_{j+1}]^{\prime})=\text{rank}([B_{j},B_{j+1}]).$
@@ 3 Types of breakpoints
Next, we show that there are two types of breakpoints.
@@
If $r_{j,j+1}>\max\{r_{j},r_{j+1}\}$, the dimension of the factor space over the combined regime $k_{j-1}^{0}+1,k_{j+1}^{0}]$ increases due to the structural break. This occurs when the factor loading matrices from regime $k_{j-1}^{0}+1,k_{j}^{0}]$ and regime $k_{j}^{0}+1,k_{j+1}^{0}]$ jointly span a larger space than either individual matrix. This space augmentation is analogous to a regression where a break in coefficients increases the number of effective regressors; specifically, the interaction of a break dummy with original regressors leads to an expanded regressor set.
Example 1
Consider a three-regime design where each regime $j$ has $r_{j}=3$ effective factors and a loading matrix $\breve{\Lambda}_{j}$. Assume that the column spaces of $\{\breve{\Lambda}_{j}\}_{j=1}^{3}$ are linearly independent. To see how our framework accommodates this setting, we set $r=9$ and $\Lambda=[\breve{\Lambda}_{1},\breve{\Lambda}_{2},\breve{\Lambda}_{3}]$ and define $B_{j}$ as a $9\times 9$ selection matrix with $I_{3}$ as the $j$-th diagonal block and zeros elsewhere. Thus, $\{\Lambda B_{j}\}_{j=1}^{3}$ collect the first, middle, and last three columns of $\Lambda$. Under this setup, any two adjacent regimes jointly span a six-dimensional factor space, implying $r_{j,j+1}=6>\max\{r_{j},r_{j+1}\}=3$ for $j=1,2$, so both breakpoints are Type 1 (singular) changes.
Example 2
Consider a scenario with $r=3$ pseudo-factors and loading matrix $\Lambda_{N\times 3}$. Regime $j$ has three effective factors, and in regime $j+1$ one factor becomes inactive after $k_{j}^{0}$, reducing the effective rank to $r_{j+1}=2$. Our framework in (3) accommodates this case by setting $B_{j}=I_{3}$ and $B_{j+1}=\text{diag}(1,1,0)$. Since the column space of $\Lambda B_{j+1}$ is a proper subspace of that of $\Lambda B_{j}$, we have $r_{j,j+1}=3=\max\{r_{j},r_{j+1}\}>\min\{r_{j},r_{j+1}\}=2$, which identifies the breakpoint as a Type 1 (singular) change.
Type 2. $r_{j,j+1}=\min\{r_{j},r_{j+1}\}$. In this case, the number of pseudo-factors in the combined regime $k_{j-1}^{0}+1,k_{j+1}^{0}]$ is the same as the number of effective factors in each regime, and we refer to this as a rotational change. Specifically, there exists a nonsingular $r\times r$ matrix $\bf R$ such that
@@
which satisfies $r_{j,j+1}=r_{j}=r_{j+1}=rank(B_{j})$. When there is one break ($m_{0}=1$) and $B_{1}$ is nonsingular, it is analogous to the rotational change for the single break setup.
Example 3
Consider $r=3$ pseudo-factors where the loading matrix in regime $j+1$ is exactly twice that of regime $j$. Our framework accommodates this by setting $B_{j}=I_{3}$ and $B_{j+1}=2I_{3}$, such that $\Lambda B_{j+1}=2\Lambda B_{j}$. Because this nonsingular scaling preserves the column space, both regimes span the same three-dimensional factor space. Hence, $r_{j,j+1}=r_{j}=r_{j+1}=3$, identifying the breakpoint as Type 2 (rotational). More generally, the same conclusion holds if $B_{j+1}=B_{j}\textbf{R}$ for a nonsingular $3\times 3$ matrix R, where the break can equivalently be interpreted as a change in the factor covariance matrix with an unchanged loading space.
@@
These two types of breaks have distinct economic interpretations. Rotational (Type 2) breaks can be interpreted as observationally equivalent to a change in the second moment of the existing factors, while leaving the loading space unchanged. In contrast, singular (Type 1) breaks represent “heterogeneous” shifts where the column space of the loading matrix changes. This corresponds to more profound structural transformations, including the emergence of new factors or the disappearance of existing ones.
@@
Notably, the analysis of multiple breakpoints is considerably more complicated than for a single breakpoint. In a multi-breakpoint setting, $r_{j,j+1}$ can be less than the total number of pseudo-factors $r$. Specifically, it is possible that both $B_{j}$ and $B_{j+1}$ are singular, yet the $j$-th breakpoint is rotational. Consequently, relying solely on the singularity of $B_{j}$ (like the DBH rule) is insufficient for determining the breakpoint type. To understand this phenomenon, consider Equations (19) and (26): when $B_{j}$ is singular (and consequently $B_{j+1}$ is singular), the pseudo-factors in regimes $j$ and $j+1$ (i.e., $F_{j}B_{j}^{\prime}$ and $F_{j+1}B_{j+1}^{\prime}$) are not of full column rank. However, if we focus solely on the data in the combined interval $k_{j-1}^{0}+1,k_{j+1}^{0}]$ and fit a factor model for observations within this interval, the pseudo-factors in this model are of full column rank $r_{j,j+1}$. In contrast, when considering the full sample for t=1,,Tt=1,...,T, the pseudo-factors in the interval $k_{j-1}^{0}+1,k_{j+1}^{0}]$ become rank-deficient. Therefore, this singularity is not directly attributable to the $j$-th breakpoint itself but rather results from the expansion of the factor loading space caused by other breakpoints, which renders the pseudo-factors in $k_{j-1}^{0}+1,k_{j+1}^{0}]$ rank-deficient. This highlights that for rotational changes, the singularity of $B_{j}$ and $B_{j+1}$ is not introduced by the $j$-th breakpoint but rather by the cumulative effect of structural changes at multiple locations. Therefore, relying solely on the singularity of $B_{j}$ and $B_{j+1}$ is insufficient to distinguish between Type 1 and Type 2 changes at the $j$-th breakpoint, as both types may be associated with singular $B_{j}$ and $B_{j+1}$. Situations involving rank deficiency for regimes before and after a rotational change occur only when multiple breakpoints are present; such scenarios do not arise in single-breakpoint cases, highlighting the complexity introduced by the multi-breakpoint setting.
The preceding discussion naturally raises the question of how to distinguish between the two types of changes in multi-breakpoint settings, particularly when the singularity of $B_{j}$ and $B_{j+1}$ (the DBH rule in a single-break setting) is no longer sufficient for distinction. The answer is to use the definition $r_{j}=r_{j+1}=r_{j,j+1}$ to determine whether the $j$-th breakpoint is rotational. Evidently, if Equation (19) holds, then $r_{j}=r_{j+1}=r_{j,j+1}$ must also hold, implying that the $j$-th breakpoint is rotational. Conversely, another natural question arises: does $r_{j}=r_{j+1}=r_{j,j+1}$ always imply the existence of a nonsingular $R$ matrix satisfying Equation (19)? This is affirmed by Proposition 1, which demonstrates that Equation (19) is a necessary and sufficient condition for defining a Type 2 break. This proposition also explains why Type 2 is named a “rotational change.”
@@
Proposition 1 can equivalently be written as follows: There does not exist a nonsingular $\bf R$} such that $B_{j}{\bf R}=B_{j+1}$ $\Leftrightarrow$ $r_{j,j+1}>\min(r_{j},r_{j+1})$, which corresponds to a type 1 change. To further illustrate this, consider two distinct cases for a setting involving two breakpoints. Case 1 involves three regimes characterized by $B_{1}$, $B_{2}$, and $B_{3}$; Case 2 involves $B_{1}$, $\tilde{B}_{2}$, and $B_{3}$. These matrices are defined as follows:
@@
However, such nonsingular $\bf R$} does not exist for $\tilde{B}_{2}$ because the second row of the product $B_{1}{\bf R}$ is always zeros, which is not equal to that of $\tilde{B}_{2}$, i.e., for any $\bf R$},
@@
For Type 1 breakpoints, we will show that the QML estimators converge to the true breakpoints with probability approaching one. In contrast, for Type 2 breakpoints, the difference between QML estimators and the true change points is stochastically bounded. The specific asymptotic theory is detailed in Section 4.
@@ 4.1 Assumptions
(i) $E\left\|f_{t}\right\|^{4}<M<\infty$, $E(f_{t}f_{t}^{{}^{\prime}})=\Sigma_{F}$, where $M$ is a positive constant, $\Sigma_{F}$ is positive definite, and $\frac{1}{k_{j}^{0}-k_{j-1}^{0}}\sum_{t=k_{j-1}^{0}+1}^{k_{j}^{0}}f_{t}f_{t}^{{}^{\prime}}\xrightarrow{p}\Sigma_{F}$ as $T\to\infty$ for $j=1,\cdots,m_{0}+1$;
@@
(iii) $B_{1},\cdots,B_{m_{0}+1}]$ is of full row rank.
@@
$\left\|\lambda_{i,j}\right\|\leq\bar{\lambda}<\infty$ for $j=1,\cdots,m_{0}+1$, $i=1,\cdots,N$, and $\left\|\frac{1}{N}\Lambda^{{}^{\prime}}\Lambda-\Sigma_{\Lambda}\right\|\rightarrow 0$ as $N\to\infty$ for some $r\times r$ positive definite matrix $\Sigma_{\Lambda}$.
@@
    For every $(i,j)$ and $\ell=1,\cdots,m_{0}+1$, $E\left|(k_{\ell}^{0}-k_{\ell-1}^{0})^{-1/2}\sum_{t=k_{\ell-1}^{0}+1}^{k_{\ell}^{0}}(e_{it}e_{jt}-E[e_{it}e_{jt}])\right|^{4}\leq M$.
@@
Define $\epsilon_{t}=f_{t}f_{t}^{{}^{\prime}}-\Sigma_{F}$. The Hájek–Rényi inequality applies to the processes $\{\epsilon_{t},t=k_{j-1}^{0}+1,\cdots,k_{j}^{0}\}$ and $\{\epsilon_{t},t=k_{j}^{0},\cdots,k_{j-1}^{0}+1\}$ for $j=1,\cdots,m_{0}+1$.
@@
For a fixed $s$ and varying $k$ (or vice-versa), there exists a constant $M<\infty$ such that for all $k-s>R\equiv(\log T)^{2}$:
(i) For each $\ell=1,\dots,T$:
$\max_{k-s>R}\frac{1}{k-s}\sum_{t=s+1}^{k}\left|\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\left[e_{i\ell}e_{it}-E(e_{i\ell}e_{it})\right]\right|^{2}=O_{p}(1);$
(ii) $\max_{k-s>R}\frac{1}{k-s}\sum_{t=s+1}^{k}\left\|\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\lambda_{i}e_{it}\right\|^{2}=O_{p}(1);$
(iii) $\max_{k-s>R}\frac{1}{k-s}\sum_{t=s+1}^{k}\left\|\frac{1}{\sqrt{NT}}\sum_{h=1}^{T}\sum_{i=1}^{N}f_{h}[e_{ih}e_{it}-E(e_{ih}e_{it})]\right\|^{2}=O_{p}(1).$